OpenStudy (anonymous):

(a^2)^1/2 how do i do problems like this?

OpenStudy (zzr0ck3r):

\((a^b)^c=a^{bc}\)

OpenStudy (anonymous):

just move it around?

OpenStudy (zzr0ck3r):

\((a^2)^{\frac{1}{2}}=a^{2*\frac{1}{2}}=a^1=a\)

OpenStudy (javk):

no, multiply the powers

OpenStudy (zzr0ck3r):

this is true for \(a\ge 0\)

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

using some logic about what you know about squares, what do we get if \(a\) is negative?